joses gas tank is 1/9 full after he buys 6 gallons it is 1/3 full how many can joses tank hold
Answer by Theo(13342)

x is the amount of gas that the tank can hold.
when he starts he has 1/9 * x gallons of gas in the tank.
after he buys 6 gallons he has 1/3 * x gallons of gas in the tank.
your equation is:
1/9 * x + 6 = 1/3 * x
subtract 1/9 * x from both sides of this equation to get:
6 = 1/3 * x - 1/9 * x
since 1/3 is the same as 3/9, your equation becomes:
6 = 3/9 * x - 1/9 * x
simplify this by performing the subtraction to get:
6 = 2/9 * x
multiply both sides of this equation by (9/2) to get:
9/2 * 6 = 9/2 * 2/9 * x
this becomes:
9/2 * 6 = x
perform the indicated operations to get:
54/2 = x which results in:
x = 27
the tank holds 27 gallons.
1/9 * 27 is equal to 3 gallons.
that's what was in the tank when you started.
add 6 gallons to this to get 9 gallons.
you have 9 gallons after you added the 6 gallons.
since the full tank is 27 gallons, the ratio of gas in the tank is now 9/27 which is the same as 1/3.